# Multi-fidelity Gaussian process surrogate modeling for regression problems in physics

### v1

@@ 4.1 Academic Examples
Before applying the problem to real-world scenarios, we test different multi-fidelity surrogate modeling methods on some academic problems shown in Table 1, each posing a different modeling challenge. We deal with two fidelity cases and train the multi-fidelity GP surrogate using fifty low-fidelity and eight high-fidelity function evaluations. The high number of low-fidelity train data ensures low predictive variance for the low-fidelity Gaussian process. We use twenty-five and eight induction points in each level for all the cases involving deep Gaussian processes. Additionally, we also train a GP on high-fidelity data without multi-fidelity features. We then compare the multi-fidelity methods to the single-fidelity GP surrogate.
@@ 4.2 Terramechanical example
As data sources for lower-fidelity models for this experiment, we used two types of models, both of which were developed at the Terramechanical lab at German Aerospace Center (DLR): TerRA [55] and SCM [53]. As a high-fidelity data source, DLR’s testbed TROLL was used [56]. This testbed was initially created to test and validate simulations for the wheel-soil interactions for extraterrestrial rover projects like MMX [57] and Scout [58]. The 44 high-fidelity runs from TROLL were also replicated in the simulation environment to achieve both higher- and lower-fidelity observations on the same set of inputs. These runs included a variety of different steering and titling angles, different rotation and horizontal velocities, and different movement trajectories and accelerations. Movement scenarios for dataset creation were described and justified in the previous paper [54]. The scheme of the overall setup of experiments is depicted in Figure 7(a).
@@
We tested all MF models discussed earlier on the terramechanical data. MF models shown here were trained with both SCM and TerRA surrogates being the lower-fidelity function and compared with each other. Table 4 shows the performance of different MF methods w.r.t. the number of high-fidelity data points used during the modeling and numerical simulation which generated the data for the lower-fidelity level. The performance of the MF models with the TerRA as a lower-fidelity data source shows marginally worse results than with SCM, however still very compatible. This is a positive result, given that the TerRA model is much simpler in implementation and less expensive in exploitation. Empirically, there is no advantage in building more than two layers of fidelity for regression tasks, if we can already approximate the medium-fidelity surrogate with high accuracy. And given an abundance of data on medium-fidelity levels, we can skip this step. Changing the number of high-fidelity points can give a hint of the lower acceptable bar when we construct the MF model.
@@
Figure 9: Performance comparison of predictions from the conventional numerical simulation methods and MF ML model. The multi-fidelity approach is computationally efficient while maintaining a high level of precision.
@@
Another advantage of the Gaussian process is the implicit handling of uncertainties. Due to its Bayesian nature, we can propagate uncertainties easily using the equation 5. It is an important part of predicting complex simulation systems, both for operational and research purposes. This adds more depth to the understanding and explainability of predictions and makes ML black boxes more transparent.
As we can see in Figure 8, the main disadvantage of the NARDGP multi-fidelity prediction is that confidence intervals are too narrow, as discussed in Section 2.3. This is a critical flaw when it comes to deploying ML models in practice, especially for highly sensitive operations like extraterrestrial rovers, where the cost of error is very high due to the inability to quickly compensate for the wrong move or repair the rover. To solve this problem, we applied several calibration methods, mentioned in Section 2.3. Isotonic regression is an algorithm enabling quantile calibration, inspired by a Platt scaling algorithm from classification [38]. Beta calibration was chosen as an algorithm for achieving distribution calibration [37]. Normal calibration [36] introduces a scaling parameter for the predicted variance, preserving a Gaussian nature of the posterior prediction. A comparison of all calibration techniques on all of the test runs can be seen in Table 5. As an example, we took NARDGP trained with 385 high-fidelity data points and calibrated it with the calibration mentioned above techniques. From this table, we can see that isotonic regression performs the best calibration, despite that it is only a global calibration of the marginal distribution. This means that the NARDGP model was constantly making overconfident predictions and global scaling of the variance solves the problem. These findings are consistent with the previous study, researching Gaussian process regression calibration for terramechanical data, but conducted only in the context of single-level medium-fidelity SCM data [61]. An example of a calibration on one test run could be seen in Figure 10, where the increased reliability of the predicted uncertainties is noticeable.
@@ Acknowledgments
Research by Vladyslav Fediukov, Michael Bergmann and Kislaya Ravi is funded by the Helmholtz Association under the ”Munich School for Data Science - MuDS”. Felix Dietrich acknowledges funding by the DFG project no. 468830823, and also association to DFG-SPP-229.

### v2

@@ 4.1 Academic Examples
Before applying the problem to real-world scenarios, we test different multi-fidelity surrogate modeling methods on some academic problems shown in Table 1, each posing a different modeling challenge. Researchers addressed the benchmarking of the surrogate models and there exists a variety of them [57]. Still, our low-fidelity function for all the test cases is the same because we want to test the capability of different methods to learn the relationship between low and high-fidelity models. We deal with two fidelity cases and train the multi-fidelity GP surrogate using fifty low-fidelity and eight high-fidelity function evaluations. The high number of low-fidelity train data ensures low predictive variance for the low-fidelity Gaussian process. We use twenty-five and eight induction points in each level for all the cases involving deep Gaussian processes. Additionally, we also train a GP on high-fidelity data without multi-fidelity features. We then compare the multi-fidelity methods to the single-fidelity GP surrogate.
@@
(a) Target functions.
(b) Model prediction for $f_{3}$ using NARGP.
Figure 7: An example of multi-fidelity gaussian process surrogates for three fidelity case. We use NARGP to create surrogate because the transformation from first to the second fidelity is non-linear.
As indicated in Section 3, we can extend the surrogate modeling to encompass more than two fidelities. As an example, we consider a three-fidelity case defined as
$f_{1}(x)=\sin\left(8\pi x\right)$
$f_{2}(x)=\sin^{2}\left(8\pi x\right)$ (29)
$f_{3}(x)=\sin^{2}\left(8\pi x\right)+x^{2}$
Figure 7(a) shows the target function for the three fidelity cases as described in Equation 4.1. The transition from the first fidelity to the second fidelity is non-linear, necessitating the use of a non-linear method to construct a surrogate. In this example, we present the results achieved using NARGP. We begin with forty, twelve, and eight randomly sampled points for each fidelity level and then execute the adaptivity algorithm for five steps. We have already observed from the results of two-fidelity cases that the posterior variance for the second fidelity under the same scenario using the NARGP model is very small. Consequently, we can utilize the posterior mean of the surrogate of the second level as the input for the third layer. Figure 7(b) shows that the surrogate aligns closely with the target function. Analogous to the two-fidelity cases, familiarity with the physical model will guide our selection of the appropriate method.
@@ 4.2 Terramechanical example
As data sources for lower-fidelity models for this experiment, we used two types of models, both of which were developed at the Terramechanical lab at German Aerospace Center (DLR): TerRA [62] and SCM [60]. As a high-fidelity data source, DLR’s testbed TROLL was used [63]. This testbed was initially created to test and validate simulations for the wheel-soil interactions for extraterrestrial rover projects like MMX [64] and Scout [65]. The 44 high-fidelity runs from TROLL were also replicated in the simulation environment to achieve both higher- and lower-fidelity observations on the same set of inputs [^2] The data is not publicly available but is available upon request.. These runs included a variety of different steering and titling angles, different rotation and horizontal velocities, and different movement trajectories and accelerations. Movement scenarios for dataset creation were described and justified in the previous paper [61]. The scheme of the overall setup of experiments is depicted in Figure 8(a).
@@
We tested all MF models discussed earlier on the terramechanical data. MF models shown here were trained with both SCM and TerRA surrogates being the lower-fidelity function and compared with each other. Table 4 shows the performance of different MF methods w.r.t. the number of high-fidelity data points used during the modeling and numerical simulation which generated the data for the lower-fidelity level. The performance of the MF models with the TerRA as a lower-fidelity data source shows marginally worse results than with SCM, however still very compatible. This is a positive result, given that the TerRA model is much simpler in implementation and less expensive in exploitation. Changing the number of high-fidelity points can give a hint of the lower acceptable bar when we construct the MF model.
@@
Figure 10: Visual comparison of the two- and three-level NARDGP surrogate performance against the test runs from the high-fidelity TROLL experiments. The results are not identical, but the performance increase is not promising, therefore we put aside proper testing of three-level MF ML surrogates for terramechanical data.
Empirically, there is no advantage in building more than two layers of fidelity for this regression task, as SCM simulation covers all the abilities of TerRA simulation and gives much more reliable results [60]. We conducted a comparison of the two-level NARDGP model performance where SCM data serves as a lower-fidelity and TROLL as the high-fidelity data source and three-level NARDGP, with new lower-fidelity level from TerRA data, so that SCM simulations become medium-level. For both cases, we used 43 high-fidelity data points from TROLL experiments, as we are first of all interested in the model’s performance with as few high-fidelity points as possible, and 1500 data points from SCM simulation. For the three-level model, we sampled 3000 data points from the TerRA simulation. Two of the examples are depicted in Figure 10 and they show a very limited difference between the performance of the two approaches. Given the limited scale of improvements we didn’t conduct a full analysis as in Table 4, but we will leave this for further work.
Figure 11: Mean squared error comparison of predictions from the conventional numerical simulation methods and MF ML model. The multi-fidelity approach is computationally efficient while maintaining a high level of precision.
@@
Another advantage of the Gaussian process is the implicit handling of uncertainties. Due to its Bayesian nature, we can propagate uncertainties easily using the equation (5). It is an important part of predicting complex simulation systems, both for operational and research purposes. This adds more depth to the understanding and explainability of predictions and makes ML black boxes more transparent.
As we can see in Figure 9, the main disadvantage of the NARDGP multi-fidelity prediction is that confidence intervals are too narrow, as discussed in Section 2.3. This is a critical flaw when it comes to deploying ML models in practice, especially for highly sensitive operations like extraterrestrial rovers, where the cost of error is very high due to the inability to quickly compensate for the wrong move or repair the rover. To solve this problem, we applied several calibration methods, mentioned in Section 2.3. Isotonic regression is an algorithm enabling quantile calibration, inspired by a Platt scaling algorithm from classification [37]. Beta calibration was chosen as an algorithm for achieving distribution calibration [36]. Normal calibration [35] introduces a scaling parameter for the predicted variance, preserving a Gaussian nature of the posterior prediction. A comparison of all calibration techniques on all of the test runs can be seen in Table 5. As an example, we took NARDGP trained with 41 high-fidelity data points and calibrated it with the calibration mentioned above techniques. From this table, we can see that isotonic regression performs the best calibration, despite that it is only a global calibration of the marginal distribution. This means that the NARDGP model was constantly making overconfident predictions and global scaling of the variance solves the problem. These findings are consistent with the previous study, researching Gaussian process regression calibration for terramechanical data, but conducted only in the context of single-level medium-fidelity SCM data [68]. An example of a calibration on one test run could be seen in Figure 12, where prediction became less overconfident and 95% internal captures almost the entire actual observed traction force signal.
@@ Acknowledgments
Research by Vladyslav Fediukov, Michael Bergmann and Kislaya Ravi is funded by the Helmholtz Association under the ”Munich School for Data Science - MuDS”(HIDSS-006). Felix Dietrich acknowledges funding by the DFG project no. 468830823, and also association to DFG-SPP-229.